OpenStudy (anonymous):

Okay can you please check what I have so far and help me with my last question?

OpenStudy (anonymous):

1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2x + 3y = 1200 2x-1200=-3y 1/3*(2x-1200)=1/3*-3y 2/3x-400=-y -2/3x+400=y 2. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. I would put a dot on (0,400) After that I would have the slope do the rest. OR Have it go 2 points to the right and 3 to the down. (To form a triangle) Then it should be something like 3,398. 3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. F(x) = 2x + 1200 4. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts.

OpenStudy (anonymous):

I need help with 4 I dont really know what to do @ganeshie8

OpenStudy (igreen):

#1 is correct..but you should write it like y = -2/3x + 400 instead..

OpenStudy (anonymous):

Okay ill change that what about 2?

OpenStudy (igreen):

For #2 you can say that you will find the y and x-intercepts and draw a line through them.

OpenStudy (anonymous):

That all i have to say?

OpenStudy (igreen):

Yep, basically just any two points on the line and draw a straight line through them.

OpenStudy (anonymous):

okay and What about 3?

OpenStudy (igreen):

Function notation is just replacing y with f(x). So your equation was y = -2/3x + 400, you're gonna change it to f(x) = -2/3x + 400.

OpenStudy (anonymous):

Okay i see. so now its the graph, How would i do this?

OpenStudy (igreen):

Use this site to graph it: https://www.desmos.com/calculator Then take a printscreen, paste it into a paint program and save it. Then put it in the worksheet.

OpenStudy (anonymous):

Okay thank you But where would the intercepts be on the graph?

OpenStudy (igreen):

You already have the y-intercept, it's 400. To find the x-intercept input 0 for y in the equation.

OpenStudy (anonymous):

So would the X intercept be 600?

OpenStudy (anonymous):

@iGreen I need help with this last question please 5. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how

OpenStudy (igreen):

Hold on.

OpenStudy (anonymous):

Okay